OpenStudy (anonymous):

How many factors does 1000 have?

OpenStudy (anonymous):

How many do you think

OpenStudy (alexandervonhumboldt2):

1000/2=500 500/2=250 250/2=125 125/5 =25 25/5=5 5/5=1

OpenStudy (anonymous):

1,2,4,5,8,10,20,25,40,50,100,125,200,250,500,1000,

OpenStudy (alexandervonhumboldt2):

factors are: 1,2,4,5,8,10,20,25,40,50,100,125,200,250,500,1000,

OpenStudy (alexandervonhumboldt2):

prime factorizatin is 2*2*2*5*5*5